# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v1

@@ 3.5 Confidence intervals
We construct two families of confidence intervals for $P_{ij}$: (i) finite‑sample, variance‑adaptive bands based on Lemma 3.1, and (ii) asymptotic normal‑approximation bands based on Theorem 3.5 together with a bias cushion from Lemma 3.2.
3.5.1 Finite‑sample Bousquet bands
Estimate $V_{ij}$ by the plug‑in
$\widehat{V}_{ij}:=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\widehat{P}_{kj}(1-\widehat{P}_{kj}),$
where $\widehat{P}_{kj}$ is the symmetrized estimator (2.4). For $\alpha\in(0,1)$ define the half‑width
$w_{ij}^{(B)}(\alpha):=\sqrt{2\widehat{V}_{ij}\log(2/\alpha)}+\frac{1}{3h}\log(2/\alpha),$ (3.6)
and the interval
$\mathrm{CI}_{ij}^{(B)}(\alpha):=[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\ \tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)]\cap[0,1].$ (3.7)
Conditionally on $(\mathcal{N}_{i}^{(-j)},P)$, Lemma 3.1 implies $\mathrm{CI}_{ij}^{(B)}(\alpha)$ attains coverage at least $1-\alpha$, up to the error from estimating $V_{ij}$ by $\widehat{V}_{ij}$.
@@ 3.5.2 Normal‑approximation intervals
The Bousquet bands (3.7) are finite‑sample and variance‑adaptive but can be conservative. The normal bands (3.9) are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
@@
1:Sample size $n$, neighborhood size $h$, trimming quantile $\tau$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
@@
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.2), and obtain $\bar{d}_{\tau}$ via (2.3)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
@@ 4 Honest tuning via cross-validation
The leave‑one‑out (LOO) construction admits a simple and honest cross‑validation (CV) scheme for selecting the tuning parameters $(h,\tau)$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held‑out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h,\tau)$. Consequently the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard concentration bounds.
4.1 Cross‑validation procedure and notation
Fix a row index $i$. Split the column indices $\{1,\dots,n\}$ into a training set $S_{i}$ and a validation set $T_{i}$ with $|T_{i}|\approx n/2$. Let $\mathcal{G}$ be a finite grid of candidate tuning pairs $(h,\tau)$. For each $(h,\tau)\in\mathcal{G}$ and each held‑out column $j\in T_{i}$ form the LOO adjacency matrix using only columns in $S_{i}$ (so that neighborhoods are constructed without using the held‑out columns) and compute the corresponding LOO predictor
$\tilde{P}_{ij}^{(-j)}(h,\tau):=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h,\tau)}A_{kj},\qquad j\in T_{i}.$
@@ 4.1 Cross‑validation procedure and notation
The empirical CV score at row $i$ is the average squared prediction error on the validation set:
$\widehat{R}_{i}(h,\tau):=\frac{1}{|T_{i}|}\sum_{j\in T_{i}}\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h,\tau)\bigr)^{2}.$ (4.1)
@@
The corresponding (validation) prediction risk is the conditional expectation of the same quantity:
@@
$R_{i}(h,\tau):=\frac{1}{|T_{i}|}\sum_{j\in T_{i}}\mathbb{E}\bigl[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h,\tau)\bigr)^{2}\bigr].$ (4.2)
Select any minimizer of the empirical score over the grid,
$(\hat{h},\hat{\tau})\in\arg\min_{(h,\tau)\in\mathcal{G}}\widehat{R}_{i}(h,\tau).$
4.2 Unbiasedness and concentration
The next lemma records unbiasedness of the CV score and its concentration properties. The proof is a straightforward application of conditional independence (Proposition 2.7) and Hoeffding/Bernstein type bounds; we give a short proof sketch and defer full details to the Appendix.
Lemma 4.1 (CV unbiasedness and concentration).
Suppose the LOO construction of Section 2 is used and Assumptions 2.1-2.2 hold. Fix $(h,\tau)$.
  1. (a)
    (Unbiasedness) $\mathbb{E}[\widehat{R}_{i}(h,\tau)]=R_{i}(h,\tau)$.
  2. (b)
    (Pointwise concentration) There exists a universal constant $C>0$ such that for any $\delta\in(0,1)$,
    $\mathbb{P}\Big(\big|\widehat{R}_{i}(h,\tau)-R_{i}(h,\tau)\big|\geq C\sqrt{\frac{\log(1/\delta)}{|T_{i}|}}\Big)\leq\delta.$
  3. (c)
    Moreover, if $|\mathcal{G}|=O(n^{c})$ for some $c>0$, then there exists $C^{\prime}>0$ such that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
    $\sup_{(h,\tau)\in\mathcal{G}}\big|\widehat{R}_{i}(h,\tau)-R_{i}(h,\tau)\big|\leq C^{\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$ (4.3)
Sketch of proof.
For fixed $(h,\tau)$ and $j\in T_{i}$, conditional on $\mathcal{N}_{i}^{(-j)}(h,\tau)$ and $P$ the predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is measurable with respect to $\{A_{k\ell}:k,\ell\neq j\}$ and hence independent of $A_{ij}$ by Proposition 2.7. This gives $\mathbb{E}[(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}\mid\mathcal{N}_{i}^{(-j)},P]$ equal to the conditional risk, and averaging over $j\in T_{i}$ yields unbiasedness. Pointwise concentration follows by applying Hoeffding’s (or Bernstein’s) inequality to the bounded summands $(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}$ and then integrating out the conditioning. The uniform bound (4.3) is obtained by a union bound over the finite grid $\mathcal{G}$. Full details are given in Appendix C. ∎
When evaluating a candidate pair $(h,\tau)$ on a validation index $j\in T_{i}$, the LOO predictor $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is constructed from the matrix $A^{(-j)}$ in which the $j$th row and column have been zeroed. Hence $\tilde{P}_{ij}^{(-j)}(h,\tau)$ is measurable with respect to $\sigma(\{A_{k\ell}:k\neq j,\ell\neq j\})$ and is therefore independent of $A_{ij}$ conditional on $P$. This ensures that the squared prediction error $(A_{ij}-\tilde{P}_{ij}^{(-j)})^{2}$ is an unbiased estimate of the conditional risk. The construction is repeated separately for each $j\in T_{i}$, so that no validation edge is ever used in forming its own predictor.
4.3 Oracle inequality for the selected tuning parameters
[…]
Theorem 4.2 (Oracle inequality).
[…]
$R_{i}(\hat{h},\hat{\tau})\leq\inf_{(h,\tau)\in\mathcal{G}}R_{i}(h,\tau)+C^{\prime\prime}\sqrt{\frac{\log(|\mathcal{G}|/\delta)}{|T_{i}|}}.$ (4.4)
Sketch of proof.
[…]
$R_{i}(\hat{h},\hat{\tau})\leq\widehat{R}_{i}(\hat{h},\hat{\tau})+\Delta\leq\widehat{R}_{i}(h^{*},\tau^{*})+\Delta\leq R_{i}(h^{*},\tau^{*})+2\Delta,$
where $\Delta=C^{\prime}\sqrt{\log(|\mathcal{G}|/\delta)/|T_{i}|}$. Rearranging yields (4.4) with $C^{\prime\prime}=2C^{\prime}$. A complete proof with constants is given in Appendix C. ∎
4.4 Practical remarks
  • Choice of $T_{i}$. A validation set of size $|T_{i}|\approx n/2$ balances estimation accuracy of the CV score and the number of training columns used to form neighborhoods; other splits (e.g., $K$-fold CV) are possible but require care to preserve the LOO independence structure.
  • Grid size. The grid $\mathcal{G}$ should be polynomial in $n$ so that the uniform deviation term in (4.4) is negligible.
  • Interpretation. The oracle inequality guarantees that the selected tuning parameters achieve near‑optimal validation risk; when combined with the inferential results of Section 3 this yields honest, data‑driven tuning without compromising coverage.
  • Tuning guidance. In practice we recommend $h\asymp c\sqrt{n\log n}$ with $c\in[1.5,3]$ as a default range and a trimming quantile $\tau\in[0.3,0.5]$. Cross-validation as described in Section 4 can be used to refine these choices for a given dataset.
@@ 5 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical Zhang-Levina-Zhu (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
@@ 5.1 Setup
$h=\lfloor 1.5\sqrt{n\log n}\rfloor,\qquad\tau=0.95.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column $j$ zeroed out, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
@@ 5.2 Evaluation metrics
We report two performance measures:
@@
For the LOO estimator we constructed:
  • Bousquet intervals:
@@
Intervals were truncated to $[0,1]$. Coverage was computed as
@@ 5.3 Results
Table 2 summarizes the results for all five graphons. In every case, the LOO smoother achieves lower MSE than the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Bousquet intervals slightly more conservative.
@@
Remark 5.1.
The block graphon exhibits negligible conditional within-block variance, so the normal-based and Bousquet intervals become narrow. Rounding the empirical coverage to three decimals yields values reported as $1.000$, although the true coverage is slightly below $1$.
5.4 Interpretation
The simulation results reveal several consistent patterns across the five graphons. First, the LOO smoother generally achieves lower row-wise MSE than the classical ZLZ smoother, with the magnitude of improvement depending on the structural complexity of the graphon. For smooth and rank-one graphons the gains are modest (around 33-$10\%$), while for block and spiky graphons the improvements are substantial (approximately 2020-$35\%$). The only exception is the highly oscillatory wiggly graphon, where the two estimators perform similarly, reflecting the intrinsic difficulty of smoothing rapidly varying structure.
Second, the LOO-based confidence intervals exhibit strong finite-sample performance. For smooth, block, and rank-one graphons, both Bousquet and Normal intervals achieve near-perfect coverage, with the Bousquet intervals slightly more conservative. Coverage decreases for the wiggly graphon (approximately $0.94$ for Bousquet and $0.89$ for Normal), which is expected given the rapid local fluctuations in the underlying graphon. Even in this challenging setting, the intervals remain reasonably accurate.
Overall, these findings support the theoretical predictions: the LOO construction restores conditional independence, stabilizes variance estimation, and improves finite-sample accuracy across a wide range of graphon structures, with the largest benefits occurring in graphons exhibiting sharp or localized features.
@@ 6 Discussion and Extensions
    Directed and weighted graphons. For directed networks, one could replace the Zhang-Levina-Zhu distance by an asymmetric distance that incorporates both in- and out-neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
@@
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree-corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.
@@
    Graphon-level estimation and functional inference. Our estimator targets the probability matrix $P$ rather than the graphon $h$ itself. An interesting direction is to feed the LOO-smoothed matrix $\tilde{P}$ into subsequent nonparametric smoothing or projection steps to obtain estimators of $h$, together with confidence bands for functionals of $h$ (such as cut functionals or graphon-based norms). The LOO structure should facilitate variance control for such higher-level functionals.
    Beyond logarithmic degrees.
    Our theoretical results were developed under a minimal expected degree of order $\log n$, which ensures that neighborhoods contain enough information for variance estimation and smoothing. The same ideas extend, at least heuristically, to more general sparsity regimes. If the minimal expected degree is $d_{n}\to\infty$, the effective number of observations contributing to each smoothed estimate is of order $\sqrt{d_{n}\log n}$, leading to a row-wise risk of order
    $R_{i}\;\lesssim\;\sqrt{\frac{\log n}{nd_{n}}}.$
    This recovers the rate $\sqrt{(\log n)/n}$ when $d_{n}\asymp\log n$, improves for denser graphs (e.g. $d_{n}\asymp n^{\beta}$), and deteriorates when $d_{n}$ grows very slowly. A full treatment of these sparse regimes would require revisiting the concentration and neighborhood geometry arguments of Zhang et al. (2017); Gao et al. (2015), and we leave such extensions to future work.
  5. 5.

### v3

@@ 3.4 Confidence intervals
Undder the canonical choice $h_{n}\asymp\sqrt{n\log n},$ we have
$h_{n}^{-1/2}=\left(\sqrt{n\log n}\right)^{-1/2}=(n\log n)^{-1/4}.$
Therefore the Berry-Esseen remainder in (3.9) satisfies
$\sup_{i,j}\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\right)-\Phi(x)\right|=O\!\big((n\log n)^{-1/4}\big).$
By Lemma 3.7, the bias component satisfies
$\sup_{i,j}|B_{ij}^{(-j)}|=O\!\big((\log n/n)^{1/4}\big).$
Hence the stochastic fluctuation scale and the bias scale are of the same order, up to logarithmic factors. In particular, to obtain asymptotically valid normal-approximation confidence intervals for $P_{ij}$, one must either undersmooth (so that the bias is $o(h_{n}^{-1/2})$) or incorporate an explicit bias correction of order $(\log n/n)^{1/4}$.
We construct two families of confidence intervals for $P_{ij}$: (i) finite-sample, variance-adaptive bands based on Lemma 3.8, and (ii) asymptotic normal-approximation bands based on Theorem 3.9 together with a bias cushion from Lemma 3.7.
3.4.1 Finite-sample Confidence Intervals via Empirical Bernstein Bounds
To construct finite-sample confidence intervals for the variance component $U_{ij}^{(-j)}$, we require a concentration inequality that depends only on observable quantities. Classical Bernstein inequalities involve the true variance $V_{ij}$, which is not directly accessible. In contrast, the Empirical Bernstein bound of Maurer and Pontil (2009) provides a deviation inequality that replaces the unknown variance with the sample variance, yielding a fully data-driven and nonasymptotic confidence interval.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is the leave-one-out estimator
$\tilde{P}_{ij}:=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}.$ (3.10)
The unbiased sample variance of the observed edges in the neighborhood is
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h_{n}(h_{n}-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}.$ (3.11)
Since the edges are binary, this expression simplifies to
$\hat{\sigma}_{ij}^{2}=\frac{h_{n}}{h_{n}-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}).$ (3.12)
We apply Theorem 4 of Maurer and Pontil (2009) to the independent bounded variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, with sample size $h_{n}$. For any $\alpha\in(0,1)$, a union bound over the upper and lower tails yields the conditional deviation inequality
$\mathbb{P}\!\left(\big|U_{ij}^{(-j)}\big|\leq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}\;\middle|\;\mathcal{N}_{i}^{(-j)},P\right)\geq 1-\alpha.$ (3.13)
We therefore define the variance-adaptive half-width
$w_{ij}^{(B)}(\alpha):=\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}.$ (3.14)
The resulting finite-sample confidence interval for the localized average is
$CI_{ij}^{(B)}(\alpha):=\Big[\tilde{P}_{ij}-w_{ij}^{(B)}(\alpha),\;\tilde{P}_{ij}+w_{ij}^{(B)}(\alpha)\Big]\cap[0,1].$ (3.15)
By construction, conditionally on $(\mathcal{N}_{i}^{(-j)},P)$, this interval captures the neighborhood average
$\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}$
with probability at least $1-\alpha$ for any finite $n$.
Remark 3.11 (Advantage of the Empirical Bound).
Unlike standard concentration inequalities that require substituting an estimated variance $\hat{V}_{ij}$ for the true variance $V_{ij}$, a step that typically degrades finite-sample guarantees and forces reliance on Slutsky’s Theorem and asymptotic normality, the Empirical Bernstein interval completely avoids the cost of studentization. The interval $CI_{ij}^{(B)}(\alpha)$ automatically adapts to the empirical variance of the neighborhood while strictly preserving its $1-\alpha$ coverage probability for any finite sample size.
@@ 3.4.2 Normal‑approximation intervals
The Empirical Bernstein bands (3.15) are finite-sample and variance-adaptive but can be conservative. The normal bands (3.17) are typically tighter but rely on asymptotic approximation and an explicit bias cushion; the two constructions provide complementary guarantees in practice.
@@
1:Sample size $n$, neighborhood size $h$, significance level $\alpha$, bias constant $c_{\mathrm{bias}}$
@@
5:For each $j$ form $A^{(-j)}$, compute distances $d_{ik}^{(-j)}$ via (2.4)
6:Build LOO neighborhoods $\mathcal{N}_{i}^{(-j)}$ of size $h_{n}$ and classical neighborhoods $\mathcal{N}_{i}$ from full AA
@@ 4 Honest tuning via cross-validation
The leave-one-out (LOO) construction admits a simple and honest cross-validation (CV) scheme for selecting the neighborhood size $h$ without invalidating subsequent inference. The key structural property is that, for indices $j$ in a held-out set, the observed edge $A_{ij}$ is independent of the corresponding LOO predictor $\tilde{P}_{ij}^{(-j)}(h)$ conditional on the LOO neighborhood $\mathcal{N}_{i}^{(-j)}(h)$. Consequently, the squared prediction error on a validation set is an unbiased estimate of the prediction risk and admits standard
4.1 Cross-validation procedure and notation
Fix a row index $i\in[n]$. Let $\mathcal{G}$ be a finite grid of candidate neighborhood sizes,
$\mathcal{G}\subset\{1,2,\dots,n-2\},$
typically chosen around the theoretical scale $h_{n}\asymp\sqrt{n\log n}$. For each $h\in\mathcal{G}$ and each held-out column $j\neq i$, form the reduced adjacency matrix $A^{(-j)}$ by deleting the $j$-th row and column. This ensures that the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed without using any information from the held-out target column $j$. Compute the corresponding leave-one-out predictor
$\tilde{P}_{ij}^{(-j)}(h):=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}(h)}A_{kj},\qquad j\neq i.$
@@ 4.1 Cross-validation procedure and notation
The empirical cross-validation score at row $i$ is the average squared prediction error across all available held-out columns:
$\widehat{R}_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}.$ (4.1)
@@
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h)$ given $P$, this evaluates to
@@
$R_{i}(h):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h)\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h),P\right].$ (4.2)
We then select any minimizer of the empirical score over the grid,
$\hat{h}\in\arg\min_{h\in\mathcal{G}}\widehat{R}_{i}(h).$
@@ 5 Simulation Study
We conducted a simulation study to evaluate the finite-sample performance of the proposed leave-one-out (LOO) neighborhood smoother and compare it with the classical (ZLZ) neighborhood smoother. All computations were implemented in Python using NumPy with a fixed random seed.
@@ 5.1 Setup
$h=\lfloor 1.5\sqrt{n\log n}\rfloor.$
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
@@ 5.2 Evaluation metrics
We report two performance measures.
@@
For the LOO estimator, we consider:
  • Empirical Bernstein intervals:
@@
Intervals are truncated to $[0,1]$. Coverage is computed as
@@ 5.3 Results
Table 2 summarizes the results for all five graphons. The LOO smoother achieves a comparable MSE to the classical ZLZ smoother. Coverage for both confidence interval constructions is close to the nominal $95\%$ level, with the Empirical Bernstein intervals slightly more conservative.
@@ 6 Discussion and Extensions
    Directed and weighted graphons. For directed networks, one could replace the Zhang et al. (2017) distance by an asymmetric distance that incorporates both in and out neighborhood information, and then apply the same LOO decoupling on the relevant columns. For weighted networks, a similar strategy could be pursued using appropriate moment-based distances built from $A$ and $A^{2}$, with boundedness or tail assumptions tailored to the edge-weight distribution.
@@
    Degree correction and heterogeneous sparsity. Many real networks exhibit highly heterogeneous degree distributions. Extending the present analysis to degree corrected graphon or blockmodel formulations would require incorporating node-specific scale factors into both the distance and the bias analysis. The LOO idea remains applicable, but the variance proxy $V_{ij}$ and the bias bounds would need to be adapted to handle varying effective sample sizes across nodes.